OpenStudy (hockeychick23):

In craps, which of the following scenarios is least likely? (Remember, for each roll you're rolling two dice, so "rolling a 7" means any combination of dice that adds to 7.) A. Roll an 8 on the first roll, and an 8 on the second roll to win the game. B. Roll a 7 on the first roll to win the game. C. Roll a 6 on the first roll, and a 6 on the second roll to win the game. D. Roll a 4 on the first roll, and a 7 on the second roll to lose the game. E. Roll a 2, 3, or 12 on the first roll to lose the game

OpenStudy (hockeychick23):

@jim_thompson5910 @ganeshie8

jimthompson5910 (jim_thompson5910):

what probabilities have you computed so far?

OpenStudy (hockeychick23):

well i got: 2: 1/36: 2.78% 3: 2/36: 5.56% 4: 3/36: 8.33% 5: 4/36: 11.10% 6: 5/36: 13.90% 7: 6/36: 16.70% 8: 5/36: 13.90% 9: 4/36: 11.10% 10: 3/36: 8.33% 11: 2/36: 5.56% 12: 1/36: 2.78%

jimthompson5910 (jim_thompson5910):

good so far

jimthompson5910 (jim_thompson5910):

let's look at "A. Roll an 8 on the first roll, and an 8 on the second roll to win the game. " what is the probability here?

OpenStudy (hockeychick23):

but I'm not sure what to do now

jimthompson5910 (jim_thompson5910):

what's the probability of rolling a single 8?

OpenStudy (hockeychick23):

5/36

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

how about 2 eights?

OpenStudy (hockeychick23):

25/1296

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

25/1296 = 0.0192 = 1.92% (roughly)

jimthompson5910 (jim_thompson5910):

"B. Roll a 7 on the first roll to win the game. " that's roughly 6/36 = 16.70% and you already got that part

jimthompson5910 (jim_thompson5910):

"C. Roll a 6 on the first roll, and a 6 on the second roll to win the game. " this is found the same way done in part A

OpenStudy (hockeychick23):

25/1296 again

jimthompson5910 (jim_thompson5910):

yep which is roughly 1.92%

jimthompson5910 (jim_thompson5910):

"D. Roll a 4 on the first roll, and a 7 on the second roll to lose the game." same idea but with different numbers per roll

OpenStudy (hockeychick23):

3/36*6/36= 18/1296

jimthompson5910 (jim_thompson5910):

yep, 18/1296 = 0.01388 = 1.388% approx

jimthompson5910 (jim_thompson5910):

"E. Roll a 2, 3, or 12 on the first roll to lose the game" be careful about the keyword "or"

OpenStudy (hockeychick23):

oh ok so what do i do different on this one?

jimthompson5910 (jim_thompson5910):

you add the individual probabilities when it comes to "or" this applies if the probabilities are mutually exclusive (they are in this case)

OpenStudy (hockeychick23):

so 4/36 or 1/9

jimthompson5910 (jim_thompson5910):

P(A or B) = P(A)+P(B) ... events A and B are mutually exclusive P(A or B or C) = P(A)+P(B)+P(C) ... events A,B,C are mutually exclusive

jimthompson5910 (jim_thompson5910):

yes, roughly 1.11%

jimthompson5910 (jim_thompson5910):

now I'm going to list out the probabilities as percentages for A through E A: 1.92% B: 16.70% C: 1.92% D: 1.388% E: 1.11%

OpenStudy (hockeychick23):

so then the answer would be E right?

jimthompson5910 (jim_thompson5910):

using percentages allows us to order the values

jimthompson5910 (jim_thompson5910):

yes, 1.11% is the smallest of the 5

OpenStudy (hockeychick23):

ok thanks!!

jimthompson5910 (jim_thompson5910):

np